OpenStudy (anonymous):

Solve. box your answer(s). x^2(3x+1)=24x

OpenStudy (anonymous):

3x^2 + x^2= 24x 4x^2=24x thats the first two steps. Distribute the x^2 and combine like terms. Try to solve from there

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

i got [3x ^{2}+x ^{2}-24x=0\]

OpenStudy (anonymous):

so from here do i set each term equal to zero?

OpenStudy (anonymous):

No u cant do that unless u have something like (x-3)(x+5)=0

OpenStudy (anonymous):

so i factor it then set them equal to zero?

OpenStudy (anonymous):

yeah u gotta factor out an x

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

do u figure this one out @daisy.paz

OpenStudy (anonymous):

not exactly. i got to \[3x ^{2}+x ^{2}-24x=0\]

OpenStudy (anonymous):

from here i don't know where to go.

OpenStudy (anonymous):

3x^3 not 3x^2

OpenStudy (anonymous):

true! thanks for catching the mistake

OpenStudy (anonymous):

yeah! way to go man @surdawi where do we go from there tho?

OpenStudy (anonymous):

factor an x

OpenStudy (anonymous):

i feel like i should factor it and then set it to zero to get the answers

OpenStudy (anonymous):

i think u should factor the x out. Then it will be in quadratic equation form

OpenStudy (anonymous):

x(3x^2+x+24)=0

OpenStudy (anonymous):

thats what i said

OpenStudy (anonymous):

okay. where did you get positive 24?

OpenStudy (anonymous):

-24 sry

OpenStudy (anonymous):

ok and from there do i factor it?

OpenStudy (anonymous):

is that [a ^{2}+b ^{2}=c ^{2}\]

OpenStudy (anonymous):

hey @daisy.paz, jbovey doesnt want me to help you with this question

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Quadratic_equation

OpenStudy (anonymous):

then plug those numbers into quadratic equation

OpenStudy (anonymous):

u get it?

OpenStudy (anonymous):

not yet.

OpenStudy (anonymous):

i have 9x^2+1x+24=0

OpenStudy (anonymous):

is that righ or wrong? i don't know where to go from here?

OpenStudy (anonymous):

no thats wrong. a=3 b=1 c=24 just plug those #'s into the quadratic equation. Read that link I posted for u

OpenStudy (anonymous):

so its [3x ^{2}+x ^{2}+24=0\]

OpenStudy (anonymous):

i mean 3x^2+x-24=0

OpenStudy (anonymous):

Thats the form u need it in to calculate what a,b, and c equal. Once you know this plug it into the quadratic equation which is this.

OpenStudy (anonymous):

c=-24 sry i keep forgetting its negative

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

lemme kno when u calculate it and ill let u kno if its right

OpenStudy (anonymous):

\[-1\pm \sqrt{288\div6}\]

OpenStudy (anonymous):

the divide by 6 should be at the bottom

OpenStudy (anonymous):

almost but it should be square root of 289 not 288. The square root of 289=17. so Its 17/6

OpenStudy (anonymous):

but the square root equals a decimal. so i think i am wrong

OpenStudy (anonymous):

-1+or- (17/6)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

pretty complicated, but once u know what a,b, and c are it should be easy.

OpenStudy (anonymous):

okay! thank you!

OpenStudy (anonymous):

wouldn't it be -1/6

OpenStudy (anonymous):

i would round 17/6 to 3. so you'll get two answers. -1+3=2 and -1-3=-4

OpenStudy (anonymous):

just let ur teacher know u rounded

OpenStudy (anonymous):

\[-\frac{ 1}{ 6 }\pm\frac{ 17}{ 6 }\] \[-\frac{ 1}{ 6 }+\frac{ 17}{ 6 }=\frac{18}{6}=3\] \[-\frac{ 1}{ 6 }-\frac{ 17}{ 6 }=\frac{16}{6}=\frac{ 8}{ 3 }\] but still missing a value

OpenStudy (anonymous):

oh yeah forgot to put -1/6. sorry bout that

OpenStudy (anonymous):

jbovey do you want to tell her the 3rd solution?

OpenStudy (anonymous):

im not exactly sure what that would be. You go ahead man, thanks for catching my mistake @surdawi

OpenStudy (anonymous):

x(3x^2+x-24)=0 -3 and 8/3 are the solutions for (3x^2+x-24) "i did an error when calculating up there" but x=0 is also a solution

OpenStudy (anonymous):

\[-\frac{ 1 }{ 6 }+\frac{ 17 }{ 6 }=\frac{ 16 }{ 6 }=\frac{ 8 }{ 3 }\] \[-\frac{ 1 }{ 6 }-\frac{ 17 }{ 6 }=-\frac{ 18 }{ 6 }=-3\]

OpenStudy (anonymous):

okay thank you so much to both of you!